# Cross-Document Event-Keyed Summarization

### v1

@@ Abstract
Event-keyed summarization (EKS) requires generating a summary about a specific event described in a document, given the document and an event representation extracted from it. In this work, we extend EKS to the cross-document setting (CDEKS), in which summaries must synthesize information from accounts of the same event given by multiple sources. We introduce SEAMuS (Summaries of Events Across Multiple Sources), a high-quality dataset for CDEKS based on an expert reannotation of the FAMuS dataset for cross-document argument extraction. We present a suite of baselines on SEAMuS, covering both smaller, fine-tuned models, as well as zero- and few-shot prompted LLMs, along with detailed ablations, and a human evaluation study—showing SEAMuS to be a valuable benchmark for this new task.
@@ 1 Introduction
Providing useful and actionable information about events requires the ability not only to extract relevant, user-specified information from documents, but also to surface that information in a readable form. Drawing on this observation, Gantt et al. (2024) recently proposed event-keyed summarization (EKS), a task that entails generating a summary of a particular event, given a document and an event representation extracted from it. EKS thus seeks to satisfy both of the above requirements—reconciling the very specific interests of information extraction (IE) end users with the more generic outputs of traditional summarization models—in order to communicate precise information about a single event in a contextualized and human-readable form. EKS can thus be understood as an event-centric form of controllable summarization (Fan et al., 2018), where the attributes to be controlled are the event and roles of interest.
@@
The details available about a given event from any one source are often incomplete, however; full understanding routinely requires synthesizing information across multiple sources. Accordingly, this work extends EKS to the cross-document setting (CDEKS), drawing on—and enhancing—the FAMuS dataset for cross-document argument extraction (CDAE) to do so (Vashishtha et al., 2024). We summarize our contributions as follows:
@@
    Based on (1), we collect and release SEAMuS, an expert-annotated dataset of single- and cross-document event-keyed summaries, and the first ever dataset for CDEKS. [^1] https://github.com/wgantt/SEAMuS
@@
    We conduct detailed ablations and a human evaluation, giving a fine-grained picture of CDEKS’s demands and of models’ current capabilities.
@@ 2 Background
(Vashishtha et al., 2024) is a dataset of short (1- to 3-sentence) English Wikipedia passages (reports) paired with much longer, genre-diverse English source documents cited by those reports. [^2] All documents are from MegaWika (Barham et al., 2023). FAMuS supports two tasks: (1) Source Validation (SV), where the goal is to determine whether a candidate source document is valid for—i.e. describes the same event as—an event identified in a provided report; and (2) Cross-Document Argument Extraction (CDAE), which entails extracting arguments for an identified event from both the report and a valid source document. Our dataset builds on the FAMuS CDAE data, which contains 1,253 report-source document pairs (split 3:1:1 across train, dev, and test), and which annotates arguments of the same target event for each document in a pair using a subset of the FrameNet ontology that is restricted to frames denoting events, states, or processes (Baker et al., 1998). A single, maximally “informative” mention is annotated for each argument, where proper names are considered more informative than nominal expressions, which in turn are more informative than pronouns (cf. Li et al., 2021). In both report and source texts, arguments are often distributed across sentences.
@@
In introducing EKS, Gantt et al. (2024) released MUCSUM, a dataset for EKS based on the classic MUC-4 template filling dataset (Sundheim, 1992). MUCSUM contains abstractive event-keyed summaries of at most three sentences for each event template in MUC-4, written so as to faithfully communicate the role of each template argument, plus any minimal additional context required for the summary to provide a standalone account of the event. Gantt et al. present baseline results on MUCSUM using fine-tuned and zero-shot prompted models, and also conduct a human evaluation of model outputs, which inspires our own human evaluation (§5).
Beyond this, there has been surprisingly little work explicitly on event-centric summarization. S Hussain et al. (2022) use event-related keywords extracted from documents to condition summarization on CCN/DailyMail (Nallapati et al., 2016) and XSUM (Narayan et al., 2018), and integrate a novel event-oriented attention mechanism into BART (Lewis et al., 2020) to obtain summaries that aim to cover all events a document reports on.
Second, Vallurupalli et al. (2022) introduce the POQue dataset, which features a wealth of annotations that aim to characterize complex events described in stories via their subevent structure and the changes undergone by individual participants. Among these annotations are process summaries, which give high-level descriptions of a complex event as a whole, as well as change summaries, which describe the changes experienced by a single participant as a result of the event.
Key differences between these works and ours are (1) our focus on single (not multiple) events (S Hussain et al., 2022); and (2) our use of an event ontology—meant to represent a user’s information need—to control summary contents (both).
@@
CDEKS is an event-centric multi-document summarization (MDS) task. Work on MDS has pursued a wide variety of goals, including synthesizing reviews (Ganesan et al., 2010; Chu and Liu, 2019, i.a.), summarizing dialogues (Chen et al., 2021; Kraaij et al., 2005, i.a.), distilling news articles (notably, via the Document Understanding [^3] https://duc.nist.gov/ and Text Analysis Conferences [^4] https://tac.nist.gov/publications/index.html; cf. Fabbri et al., 2019), and generating analytical reports (Mayfield et al., 2024).
Two lines of MDS work are especially relevant. First, the CDEKS task may be viewed as an instance of what Ma et al. (2022) call query-oriented MDS, in which a query expressing the kind of information to be summarized is provided as input alongside the documents. [^5] This contrasts with most MDS tasks, in which a summary’s focus is defined implicitly by the topical similarity of the input documents (e.g. as determined by clustering). Queries in prior work take the form of natural language statements—e.g. Wikipedia article titles (Liu and Lapata, 2019)—or questions—e.g. users’ search queries (Pasunuru et al., 2021). In contrast, our queries are structured event representations, as may be generated by an event extraction system.
Second, among MDS datasets, our SEAMuS dataset is arguably most similar to WCEP (Gholipour Ghalandari et al., 2020) and WikiSum (Liu et al., 2018) in drawing on Wikipedia passages and their source articles, but differs from these in focusing on single events (WikiSum) and in using structured event representations to guide summary generation (both). [^6] Cf. Schuster et al. (2024) for a related QA task.
Figure 2: An example from our SEAMuS dataset. Report documents (bottom left) are Wikipedia passages that describe some event (top right) and that cite a longer (non-Wikipedia) source article (bottom right) as evidence, with event arguments annotated in both documents. SEAMuS features simple summaries of these events based on only the report (top left) as well as enriched, cross-document summaries based on both the report and its source, which are typically able to recover additional information about the event (here, the Crime).
@@ 3 Annotation
Annotation of SEAMuS was divided into two phases. In the first phase, abstractive report summaries were written for each event in FAMuS (see §2) based only on its report document, and were then annotated for event arguments (§3.1). In the second phase, abstractive cross-document summaries were written for each event based jointly on its report and source documents, and were then annotated for event arguments as in the first phase (§3.2). In both phases, annotators were instructed to amend spurious, missing, or otherwise incorrect argument annotations in the report or source document before writing their summary. Thus, both phases involve (1) correcting existing FAMuS argument annotations; (2) writing a summary based on the corrected annotations; and (3) annotating arguments in the summary. The phases differ only in the documents on which the summaries are based (report only vs. report and source). All annotations were performed by authors of this work. [^7] Additional annotation details are in Appendix E.
@@ 3.1 Phase 1: Report Summaries
Similar to the summaries in MUCSUM2), the report summaries in SEAMuS are concise summaries of a single event as recounted in a single document (a FAMuS report) that aim to faithfully represent the role of each participant and to provide the minimum additional context needed to produce an accurate, standalone account of the target event. Although the FAMuS report documents are already relatively short (typically, 2-3 sentences), they often discuss multiple events. Thus, the report summaries are further distilled descriptions focused on just one event from the report.
@@
Table 1: Summary statistics for the SEAMuS report (left) and cross-document summaries (right). See Table 5 for more. Test statistics intentionally omitted.
@@ 3.2 Phase 2: Cross-Document Summaries
Five of the authors completed the Phase 2 annotation, with all summaries and arguments singly annotated as in Phase 1. Items from all three splits were randomly and evenly distributed to the five annotators. Given the complexity of the Phase 2 task, annotation was performed in two parts, using adapted versions of Vashishtha et al.’s (2024) interface for FAMuS CDAE annotation. [^8] Interface source code was obtained from Vashishtha et al.. Screenshots are shown in Appendix E.
In the first part, annotators corrected FAMuS argument annotations in the source documents and then wrote the cross-document summary based jointly on the report and source texts and their corrected arguments. Annotators were encouraged to use the most informative mention of an argument across both the report and source documents, but again were allowed to make alterations for clarity.
In the second part, annotators annotated arguments in the summaries from the first part. As in Phase 1, all annotators were provided with detailed instructions and completed a 10-example practice annotation before doing the main task. Summary argument annotations were again validated for length and to ensure that they featured as many arguments for a given role as the maximum number annotated for that role between the report and source texts.
@@ 4.1 Overview
We present experiments on both the report (§4.2) and cross-document (§4.3) summarization tasks. In the report task (single-document EKS), both the report and its annotated event are provided as input. The cross-document task (CDEKS) is analogous, but also includes the corresponding source document and its event annotation as input. These initial experiments condition on gold event structures; §4.4 then systematically evaluates the impact of degraded extractions on summary quality. Finally, in §4.5, we briefly discuss some ablations on the input inspired by similar ones from Gantt et al. (2024), with full results in Appendix F.
@@
We benchmark SEAMuS using models of two types. First, we consider several classic pretrained encoder-decoder models widely used for summarization: BART (Lewis et al., 2020), PEGASUS (Zhang et al., 2020), and T5 (Raffel et al., 2020), fine-tuning the large versions of all three on the SEAMuS training data. Second, we consider some of the latest proprietary LLMs, evaluated in both the zero- and few-shot settings: GPT-4o [^9] https://openai.com/index/hello-gpt-4o/, GPT-4o Mini (GPT-4o M in Table 2) [^10] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, Claude 3 Haiku (Claude H) [^11] https://www.anthropic.com/news/claude-3-haiku, and Claude 3.5 Sonnet (Claude S) [^12] https://www.anthropic.com/news/claude-3-5-sonnet. For the few-shot examples, we use the three examples from the train split whose frame matches that of the target example. Finally, we also give results for a report baseline (RB) that treats the report text itself as the predicted summary.
@@
Given EKS’s focus on producing summaries that recover specific pieces of information—as represented by an event’s roles—we report several other metrics that evaluate this. First, we report CEAF-REE (CR; Du et al., 2021), a form of argument $\text{F}_{1}$ that allows us to compare arguments extracted from a predicted summary against those in a reference summary, aligning arguments based on exact match. [^13] Appendix F has results with a soft variant of this metric. Following Gantt et al. (2024), we train the event extraction model of Xia et al. (2021) [^14] https://hub.docker.com/r/hltcoe/lome on SEAMuS and use it to extract arguments from the predicted summaries, constraining extraction to arguments that fill roles of the target event only.
@@
Table 2: Report and Cross-Document summarization results on SEAMuS. Best overall results are bolded; and denote best zero- and few-shot results, respectively. S=setting; RB=report baseline; ZS=zero-shot; FS=few-shot; FT=fine-tuned. See §4.1 for an explanation of metrics; higher is better for all. See Tables 8 and 9 for 95% CIs. Best A and F results ignore RB, for reasons explained in Appendix F.
@@ 4.2 Report Summarization
For the Claude and GPT models, our system prompt tasks the model with analyzing and summarizing a specific event. The user prompt provides more specific task instructions, followed by the full report text, and a description of the target event that includes (1) the frame name and definition from FrameNet; (2) the trigger; and (3) a bulleted list, where each item includes a role name, its definition, and the arguments annotated for that role. In the few-shot setting, we format the three few-shot examples (see §4.1) the same way, but with the target summary shown at the end of each. We set temperature to 0.7 and the max new tokens to 256, leaving other API defaults unchanged. [^16] Appendix C has further details on models and prompts.
@@ 4.3 Cross-Document Summarization
The setup for the cross-document task is similar to that of the report task, but adds the source text and its annotated event to the input alongside the report text and its event. As the source texts are full web articles, most are quite long (e.g. dev texts average almost 62 sentences and over 1,500 words). While this is no obstacle for the LLMs, the smaller models do not support contexts of this size. Thus, to enable a fair comparison across models, we apply a sentence retriever to the source, using the report text as a query to select the top $k$ most relevant sentences to use as context. [^17] This approach can also be justified by the fact that typically only a small portion of the source concerns the event. We consider $k\in\{3,\ldots,10\}$ and selected the maximum value such that $\geq 95\%$ of the resulting dev set contexts would fit untruncated in the input, yielding $k=7$. We experimented with the dense retrievers all-mpnet-base-v2 (based on MPNet; Song et al., 2020) and e5-large-v2 (Wang et al., 2022), but obtained our best results with BM25 (Robertson et al., 2009), which we use in all experiments. [^18] Models were evaluated on recall of annotated arguments in the retrieved contexts for the dev set for fixed $k$. At $k=7$, BM25 recovered $\sim 76\%$ of annotated source arguments.
We use the same training and inference settings from §4.2, and refer the reader to Appendices B and C for more details on input formatting.

### v2

@@ Abstract
Event-keyed summarization (EKS) requires summarizing a specific event described in a document given the document text and an event representation extracted from it. In this work, we extend EKS to the cross-document setting (CDEKS), in which summaries must synthesize information from accounts of the same event as given by multiple sources. We introduce SEAMuS (Summaries of Events Across Multiple Sources), a high-quality dataset for CDEKS based on an expert reannotation of the FAMuS dataset for cross-document argument extraction. We present a suite of baselines on SEAMuS—covering both smaller, fine-tuned models, as well as zero- and few-shot prompted LLMs—along with detailed ablations and a human evaluation study, showing SEAMuS to be a valuable benchmark for this new task.
@@ 1 Introduction
Providing useful information about events requires the ability not only to extract relevant, user-specified information from documents, but also to present that information in a readable form. Drawing on this observation, Gantt et al. (2024) recently proposed event-keyed summarization (EKS), a task that entails summarizing a particular event, given a document and an event representation extracted from it. EKS thus seeks to satisfy both requirements—reconciling the specific information needs of IE end users with the more generic outputs of traditional summarization models—in order to communicate precise information about a single event in a contextualized and readable form. EKS can thus be viewed as event-centric controllable summarization (Fan et al., 2018), where the controlled attributes are the event and roles of interest.
@@
However, adequately understanding a particular event often requires synthesizing information across multiple sources—evidenced in part by the rapidly growing interest in retrieval augmented generation (RAG; Lewis et al., 2020b). Accordingly, this work extends EKS to the cross-document setting (CDEKS), drawing on—and enhancing—the FAMuS dataset for cross-document argument extraction (CDAE) to do so (Vashishtha et al., 2024). We summarize our contributions as follows:
@@
    Based on (1), we collect and release SEAMuS, an expert-annotated dataset of single- and cross-document event-keyed summaries—the first ever dataset for CDEKS. [^1] https://github.com/wgantt/SEAMuS
@@
    We conduct fine-grained ablations and a human evaluation, detailing CDEKS demands as a task as well as models’ current capabilities.
@@ 2 Background
(Vashishtha et al., 2024) is a dataset of short English Wikipedia passages (reports) paired with much longer, genre-diverse English source documents cited by those reports. [^2] All documents are from MegaWika (Barham et al., 2023). FAMuS supports two tasks: (1) Source Validation (SV), where the goal is to determine whether a candidate source document is valid for—i.e. describes the same event as—an event identified in a provided report; and (2) Cross-Document Argument Extraction (CDAE), which entails extracting arguments for an identified event from both the report and a valid source document. SEAMuS builds on the FAMuS CDAE data, which contains 1,265 report-source document pairs (split 3:1:1 across train, dev, and test), and annotates arguments of the same target event for each document in a pair using a subset of the FrameNet ontology restricted to frames denoting events, states, or processes (Baker et al., 1998). A single, maximally “informative” mention is annotated for each argument, where proper names > nominal expressions > pronouns (see Li et al., 2021b). In both report and source texts, arguments may be distributed across sentences.
@@
In introducing EKS, Gantt et al. (2024) released MUCSUM, an EKS dataset based on the classic MUC-4 template filling dataset (Sundheim, 1992). MUCSUM contains abstractive event-keyed summaries for each event template in MUC-4, written so as to faithfully express the role of each template argument, plus any minimal additional context required for the summary to act as a standalone account of the event. Gantt et al. present baselines on MUCSUM, and also conduct a human evaluation of model outputs, which inspires our own (§5).
Other event-centric summarization research has focused on timeline summarization (TLS), which constructs chronological lists of events, often with timestamps and usually based on multiple documents (Allan et al., 2001; Chieu and Lee, 2004; Li et al., 2021a; Rajaby Faghihi et al., 2022, i.a.). Beyond TLS, S Hussain et al. (2022) use extracted event-related keywords to condition single-document summarization, and integrate an event-oriented attention mechanism into BART to encourage models to cover all events discussed. Additionally, Vallurupalli et al. (2022) introduce the POQue dataset, which has annotations that characterize the subevent structure of complex events in stories and the changes undergone by their participants. Among these annotations are process summaries, which give high-level descriptions of a complex event, and change summaries, which describe the changes experienced by a participant as a result.
@@
CDEKS is an event-centric multi-document summarization (MDS) task. Work on MDS has pursued a variety of goals, including synthesizing reviews (Ganesan et al., 2010; Chu and Liu, 2019, i.a.), summarizing dialogues (Kraaij et al., 2005; Chen et al., 2021, i.a.), distilling news articles (notably, via DUC [^3] https://duc.nist.gov/ and TAC [^4] https://tac.nist.gov/publications/index.html), and generating reports (Mayfield et al., 2024). Event-centric MDS datasets include MultiNews (Fabbri et al., 2019) and DiverseSumm (Huang et al., 2024), which focus on new stories, but SEAMuS is most similar to Auto-hMDS (Zopf, 2018) and WCEP (Gholipour Ghalandari et al., 2020) in being built on Wikipedia articles and their sources.
CDEKS departs from all of these, however, in responding to an explicit information need. It is thus an event-centric form of query-oriented MDS (Ma et al., 2020), where a query expressing the kind of information to be summarized is provided as additional input. But whereas queries from prior work are given in natural language—e.g. article titles (Liu and Lapata, 2019) or web searches (Pasunuru et al., 2021)—ours are structured event representations, drawing on the IE tradition of leveraging event ontologies to encode information needs, and enabling extraction-to-summarization pipelines.
Our Work
We summarize three key differences between prior work and our own. We focus on:
  1. 1.
    Synthesizing information about a single event across multiple sources. Both multi-event (e.g. TLS) and single-source (e.g. EKS) summarization have their place, but many practical information needs depend on the rich understanding of an individual event that is attainable only via cross-source synthesis.
  2. 2.
    Responding to a specific event-centric information need, not generically summarizing event-related content (contra S Hussain et al., 2022; Vallurupalli et al., 2022).
  3. 3.
    Leveraging rich, structured event representations to achieve (1) and (2)—not short, unstructured queries like web searches (Pasunuru et al., 2021) or topics (Allan et al., 2001; Rajaby Faghihi et al., 2022, i.a.).
Figure 2: An example from our SEAMuS dataset. Report documents (bottom left) are Wikipedia passages that describe some event (top right) and that cite a longer (non-Wikipedia) source article (bottom right) as evidence, with event arguments annotated in both documents. SEAMuS features simple summaries of these events based on only the report (top left) as well as enriched, cross-document summaries based on both the report and its source, which typically contain additional information about the event (here, the Crime). Appendix A has further examples.
@@ 3 Annotation
Annotation of SEAMuS was divided into two phases. In the first phase, abstractive report summaries were written for each event in FAMuS (see §2) based only on its report document, and were then annotated for event arguments (§3.1). In the second phase, abstractive cross-document summaries were written for each event based jointly on its report and source documents, and were then annotated for event arguments as in the first phase (§3.2). In both phases, annotators were instructed to amend spurious, missing, or otherwise incorrect argument annotations in the report or source document before writing their summary. Thus, both phases involve (1) correcting existing FAMuS argument annotations; (2) writing a summary based on the corrected annotations; and (3) annotating arguments in the summary. The phases differ only in the documents on which the summaries are based (report only vs. report and source). All annotations were performed by authors of this work. [^5] Appendix E has additional details and agreement results.
@@ 3.1 Phase 1: Report Summaries
Similar to the summaries in MUCSUM2), the report summaries in SEAMuS are concise summaries of a single event as recounted in a single document (a FAMuS report) that aim to faithfully represent the role of each participant and to provide the minimum additional context needed to serve as an accurate, standalone account of the event. Although the FAMuS report documents are already relatively short (typically, 2-3 sentences), they often discuss multiple events. [^6] E.g. for reports in the SEAMuS train split, the MegaWika dataset (Barham et al., 2023), from which the reports are taken, has an average of 21.4 FrameNet frames annotated. Thus, the report summaries are further distilled descriptions focused on just one event from the report.
@@
Table 1: Summary statistics for the SEAMuS report and cross-document summaries. See Table 7 for more.
@@ 3.2 Phase 2: Cross-Document Summaries
Five of the authors completed the Phase 2 annotation, with all summaries and arguments singly annotated as in Phase 1. Items from all three splits were randomly and evenly distributed to the five annotators. Given the complexity of the Phase 2 task, annotation was performed in two parts, using adapted versions of Vashishtha et al.’s (2024) interface for FAMuS CDAE annotation. [^7] Interface source code was obtained from Vashishtha et al. Screenshots are shown in Appendix E.
In Part A, annotators corrected FAMuS argument annotations in the source documents and then wrote the cross-document summary based jointly on the report and source texts and their corrected arguments. Annotators were encouraged to use the most informative mention of an argument across both the report and source documents, but again were allowed to make alterations for clarity.
In Part B, annotators annotated arguments in the summaries from Part A. As in Phase 1, all annotators were provided with detailed instructions and completed a 10-example practice annotation before doing the main task. Summary argument annotations were again validated for length and to ensure that they featured as many arguments for a given role as the maximum number annotated for that role between the report and source texts.
@@ 4 Experiments
Table 2: Report and Cross-Document summarization results on SEAMuS. Best overall results are bolded; and denote best zero- and few-shot results, respectively. S=setting; RB=report baseline; ZS=zero-shot; FS=few-shot; FT=fine-tuned. See §4.1 for an explanation of metrics; higher is better for all. See Tables 10 and 11 for 95% CIs. Best A and F results exclude RB, for reasons explained in Appendix F.
@@ 4.1 Overview
We present experiments on both the report (§4.2) and cross-document (§4.3) summarization tasks. In the report task (single-document EKS), both the report and its annotated event are provided as input. The cross-document task (CDEKS) is analogous, but also includes the corresponding source document and its event annotation as input. Next, in §4.4, we briefly discuss some ablations on the input inspired by similar ones from Gantt et al. (2024), with full results in Appendix F. Finally, §4.5 evaluates the impact of degraded argument extractions on summary quality.
@@
We benchmark SEAMuS using models of two types. First, we consider several classic pretrained encoder-decoder models widely used for summarization: BART (Lewis et al., 2020a), PEGASUS (Zhang et al., 2020), and T5 (Raffel et al., 2020), fine-tuning the large versions of all three on the SEAMuS training data. Second, we consider some of the latest proprietary LLMs, evaluated in both the zero- and few-shot settings: GPT-4o [^8] https://openai.com/index/hello-gpt-4o/, GPT-4o Mini (GPT-4o M in Table 2) [^9] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, Claude 3 Haiku (Claude H) [^10] https://www.anthropic.com/news/claude-3-haiku, and Claude 3.5 Sonnet (Claude S) [^11] https://www.anthropic.com/news/claude-3-5-sonnet. For the few-shot examples, we use the three examples from the train split whose frame matches that of the target example. Finally, we also give results for a report baseline (RB) that treats the report text itself as the predicted summary.
@@
Given EKS’s focus on producing summaries that recover specific pieces of information—as represented by an event’s roles—we report several other metrics that evaluate this. First, we report CEAF-REE $\text{F}_{1}$ (CR; Du et al., 2021), a form of argument $\text{F}_{1}$ that allows us to compare arguments extracted from a predicted summary against those in a reference summary, aligning arguments based on exact match. [^12] Appendix F reports a soft match variant of this metric. Following Gantt et al. (2024), we train the event extraction model of Xia et al. (2021) [^13] https://hub.docker.com/r/hltcoe/lome on SEAMuS and use it to extract arguments from the predicted summaries, constraining extraction to arguments that fill roles of the target event only.
@@ 4.2 Report Summarization
For the Claude and GPT models, our system prompt asks the model to analyze and summarize a specific event. The user prompt provides more detailed task instructions, followed by the full report text, and a description of the target event that includes (1) the frame name and definition from FrameNet; (2) the trigger; and (3) a bulleted list, where each item includes a role name, its definition, and the arguments annotated for that role. In the few-shot setting, we format the three few-shot examples (see §4.1) the same way, but with the target summary shown at the end of each. We set temperature to 0.7 and the max new tokens to 256, leaving other API defaults unchanged. [^15] Appendix C has further details on models and prompts.
@@ 4.3 Cross-Document Summarization
The setup for the cross-document task is similar to that of the report task, but adds the source text and its annotated event to the input alongside the report text and its event. As the source texts are full web articles, most are long (e.g. dev texts average almost 62 sentences and over 1,500 words). While this is no obstacle for the LLMs, the smaller models do not support contexts of this size. Thus, to enable a fair comparison across models, we apply a sentence retriever to the source, using the report text as a query to select the top $k$ most relevant sentences to use as context. [^16] This approach can also be justified by the fact that typically only a small portion of the source concerns the event. We consider $k\in\{3,\ldots,10\}$ and selected the maximum value such that $\geq 95\%$ of the resulting dev set contexts would fit untruncated in the input, yielding $k=7$. We experimented with the dense retrievers all-mpnet-base-v2 (based on MPNet; Song et al., 2020) and e5-large-v2 (Wang et al., 2022), but obtained our best results with BM25 (Robertson et al., 2009), which we use in all experiments. [^17] Models were evaluated on recall of annotated arguments in the retrieved contexts for the dev set for fixed $k$. At $k=7$, BM25 recovered $\sim 76\%$ of annotated source arguments.
We use the same training and inference settings from §4.2; see Appendices B, C for further details.